# On the Selection Stability of Stability Selection and Its Applications

### v2

@@ Abstract
Stability selection is a widely adopted resampling-based framework for high-dimensional variable selection. This paper seeks to broaden the use of an established stability estimator to evaluate the overall stability of the stability selection results, moving beyond single-variable analysis. We suggest that the stability estimator offers two advantages: it can serve as a reference to reflect the robustness of the results obtained, and help identify an optimal regularization value to improve stability. By determining this value, we calibrate key stability selection parameters, namely, the decision threshold and the expected number of falsely selected variables, within established theoretical bounds. The asymptotic distribution of the stability estimator allows us to observe convergence of stability values over successive sub-samples. This approach sheds light on the required number of sub-samples addressing a notable gap in prior studies. Pareto optimality of the proposed regularization value is also discussed. The stabplot R package is developed to facilitate the use of the plots featured in this manuscript, supporting their integration into further statistical analysis and research workflows.
Keywords: Bioinformatics, Feature Selection, Hyperparameter Tuning, Lasso, Variable Selection
@@ 1 Introduction
Stability selection, a widely recognized resampling-based variable selection framework, performs selection by examining how frequently variables are chosen across multiple randomly selected resamples (Meinshausen and Bühlmann, 2010). Meinshausen and Bühlmann (2010) analyzed the selection frequencies for individual variables over a range of regularization values, presenting the results in a plot known as the ‘stability path’. This plot facilitates the identification of variables that are consistently selected over the regularization grid. In this paper, we shift the focus from examining the stability of individual variables to evaluating the overall stability of the stability selection results. Employing an established stability estimator, we assess the stability of the results in its entirety and introduce what we call ‘Stable Stability Selection’, which identifies the smallest regularization value that yields highly stable results.
@@
The stability of a variable selection method refers to the consistency with which it selects variables across different training sets drawn from the same underlying distribution (Kalousis et al., 2007). The stability selection approach described by Meinshausen and Bühlmann (2010) uses sub-samples that are half the size of the sample size of the original dataset to perform variable selection. Stability selection enables the identification of variables that are consistently recognized as important across the majority of sub-samples given a regularization value, considering them as stable variables. Furthermore, Meinshausen and Bühlmann (2010) established an asymptotic upper-bound for the Per-Family Error Rate, which represents the expected number of falsely selected variables. Later, Shah and Samworth (2013) introduced the complementary pairs stability selection method, which computes selection frequencies by counting how often a variable is included in models fitted to complementary $50\%$ sub-samples. Shah and Samworth (2013) introduced more rigorous upper-bounds for the expected number of variables selected with low selection frequencies.
@@
There is a substantial body of research on the stability of variable selection methods, where numerous stability measures have been proposed. For more comprehensive information, we refer to Kuncheva (2007), Nogueira et al. (2018), and Sen et al. (2021). Nogueira et al. (2018) conducted an extensive literature review on the topic, consolidating the desirable properties of stability measures into five mathematical conditions. A brief overview of these properties will be provided in the following paragraph. Nogueira et al. (2018) showed that none of the stability measures previously introduced satisfied all five conditions. In response, Nogueira et al. (2018) proposed a novel stability measure, which was the first in the literature to meet all these criteria, generalizing several previous works such as in Kuncheva (2007).
Here, we give a concise review of the five conditions proposed by Nogueira et al. (2018). The first property stipulates that the stability measure should function effectively regardless of variations in the number of selected variables. The second property asserts that the stability measure should be a strictly decreasing function of the sample variance of the binary selection status of each variable across random resamples. The third property states that the stability measure should be bounded by specific constants. The fourth property implies that maximum stability is achieved when the selection results across all resamples are identical, and vice versa. Finally, the fifth property posits that if all possible selection results are equally likely to occur across random resamples, the expected value of the stability measure should remain constant.
@@
In this paper, we apply the stability estimator proposed by Nogueira et al. (2018) to the stability selection results, and proxify the stability values to find the optimal regularization value that achieves high stability with the least possible loss in prediction accuracy, if applicable. The optimal regularization value then is employed to calibrate two critical parameters of stability selection: the decision-making threshold and the expected number of falsely selected variables, balancing them with respect to one another. We also use the optimal regularization value to define a new variable selection criterion, called stable stability selection, prioritizing the selection of variables associated with highly stable results, rather than those mostly selected in their best-case scenarios over the regularization grid, which is the primary criterion in the traditional stability selection framework. In addition, the convergence value of the stability estimator serves as a reference to assess the reliability of the results obtained. The point at which convergence occurs provides valuable insight into the required number of sub-samples, an aspect for which we found no dedicated research addressing its determination.
@@
Nogueira et al. (2018) suggested that stability and prediction accuracy can be interpreted through the lens of a Pareto front Pareto (1896). We refer to this methodology as stability-accuracy selection. Building on this perspective, we demonstrate that, given two justifiable assumptions, the proposed regularization value constitutes a Pareto-optimal solution within this context.
@@
The use of stability to enhance variable selection has been a subject of interest since at least the work of Liu et al. (2010). Among the existing literature, the study by Sun et al. (2013) is perhaps the most closely related to the present work and has continued to attract interest, as demonstrated by contributions such as Yang et al. (2020), Reavie (2021), Wen et al. (2023), and Lee and Chen (2024). In Sun et al. (2013), the regularization parameter of penalized regression models is selected based on a stability-driven criterion. Nonetheless, our methodology differs from theirs in several key aspects. Most notably, the stability measure they employ, originally introduced by Cohen (1960), is based solely on the intersection of some sets and does not account for selection variance. As argued by Nogueira et al. (2018), a stability measure should decrease as selection variance increases, a property not satisfied by the measure used in Sun et al. (2013). Furthermore, our approach is embedded directly within the stability selection framework, whereas the method in Sun et al. (2013) is designed to be used externally. Another contribution of this paper, which distinguishes it from Sun et al. (2013), is the use of stability values to inform the required number of sub-samples within the stability selection framework.
@@
The structure of this paper is organized as follows. Section 2 outlines the proposed methodology. Section 3 describes the real and synthetic datasets used in the paper. The results of applying the proposed method, along with the demonstration of Pareto optimality of the proposed regularization value, are presented in Section 4.
@@ 2 Methodology
The following procedure is applied to each regularization value $\lambda\in\Lambda$. In each iteration of the stability selection process, a sub-sample, comprising half the size of the original dataset $\mathcal{D}$, is randomly selected. The Lasso model is then fitted to this sub-sample using the fixed value of $\lambda$, and the binary selection outcomes are recorded as a row in the binary selection matrix $M(\lambda)\in\operatorname{mat}(B\times p)$, where $B$ denotes the number of sub-samples. Therefore, in the end, we obtain $|\Lambda|$ distinct selection matrices $M(\lambda)$. In this context, $M(\lambda)_{bj}=1$ indicates that the $j$th variable is identified as part of the signal set $\hat{S}(\lambda)$ when Lasso is applied to the $b$th sub-sample given $\lambda$. In contrast, $M(\lambda)_{bj}=0$ signifies that the $j$th variable is classified as belonging to the noise set $\hat{N}(\lambda)$ after applying Lasso to the $b$th sub-sample given $\lambda$.
@@
where $s_{j}^{2}$ denotes the unbiased sample variance of the binary selection statuses of the $j$th variable, while $q(\lambda)$ denotes the average number of variables selected under the regularization parameter $\lambda$. The stability estimator $\hat{\Phi}(\cdot)$ is bounded by $\left[-\frac{1}{B-1},1\right]$ (Nogueira et al., 2018); the larger the $\hat{\Phi}(\cdot)$, the more stable is $M(\lambda)$. To interpret the stability values of Equation (2), Nogueira et al. (2018) adopted the guidelines proposed by Fleiss et al. (2004). According to Nogueira et al. (2018), stability values that exceed $0.75$ indicate excellent agreement between selections beyond what would be expected by random chance, while values below $0.4$ signify poor agreement between them. Based on Nogueira et al. (2018), stability values that reside between these thresholds are categorized as indicative of intermediate to good stability.
@@
The threshold value of $0.75$ is somewhat arbitrary and may not be attainable in certain situations. As we will see in Section 4, $\lambda_{\text{stable}}$ may not exist in certain practical applications due to procedural instability, which prevents the stability values from exceeding $0.75$. In such cases, we propose
@@
Both upper-bounds of the Per-Family Error Rate, mentioned in Section 1, provided by Meinshausen and Bühlmann (2010) and Shah and Samworth (2013), depend on the values of $\lambda$ and $\pi_{\text{thr}}$. The calibration of these two values requires the arbitrary selection of one of these two parameters, which can be challenging to justify (Bodinier et al., 2023). By adopting the optimal regularization value, that is $\lambda=\lambda_{\text{stable}}$ or $\lambda=\lambda_{\text{stable-1sd}}$, the upper-bound can be explicitly tailored to the data owner’s preferences. If a pre-determined $\pi_{\text{thr}}$ is chosen, the Per-Family Error Rate is deterministically obtained; alternatively, if a fixed pre-determined Per-Family Error Rate value is desired, the corresponding $\pi_{\text{thr}}$ is enforced.
@@
As outlined in Section 1, as the number of sub-samples increases, the convergence of $\hat{\Phi}(\cdot)$ to the true stability $\Phi$ is ensured. To illustrate this, we define $M^{(t)}(\lambda)$ as the matrix containing the selection results from the first $t$ sub-samples for a given $\lambda\in\Lambda$, that is, the first $t$ rows of $M(\lambda)$. The objective is to evaluate the stability of $M^{(t)}(\lambda)$ across successive sub-samples in order to determine the appropriate cut-off point for the process, that is, the number of sub-samples required to achieve convergence in stability. Beyond this threshold, additional sub-samples do not significantly change the stability of $M(\lambda)$. We propose plotting the stability values of the selection matrix over the sequential sub-sampling, that is $\hat{\Phi}(M^{(t)}(\lambda));\;t=2,3,\ldots B$, against $t$ to monitor the convergence status of the stability values. Given that the asymptotic distribution of $\hat{\Phi}(\cdot)$ is established, a confidence bound can be drawn along the curve to reflect the inherent uncertainty of the stability estimator.
In Section 4, we present the results obtained from the application of the proposed approach to synthetic and real datasets, which are introduced in Section 3.
@@ 3 Datasets
We evaluate the proposed methodology by conducting evaluations on synthetic and two real bioinformatics datasets, as detailed below.
@@
As synthetic data, we consider a dataset with a sample size of $n=50$ that includes $p=500$ predictor variables. In this scenario, the predictor variables $\boldsymbol{x}^{\intercal}_{i}$ are independently drawn from the Normal distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma\in\operatorname{mat}(p,p)$ has a diagonal of ones and $\sigma_{jk}=\rho^{|j-k|}$ for $j\neq k$ where $\rho\in\{0.2,0.5,0.8\}$. The response variable is linearly dependent solely on the first two predictor variables, characterized by the coefficient vector $\boldsymbol{\beta}=(1.5,1.1,0,\ldots 0)^{\top}$, and the error term $\boldsymbol{\varepsilon}$ is an i.i.d. sample from the standard Normal distribution $\mathcal{N}(0,1)$.
@@
For the first real example, we use the well-established ‘riboflavin’ dataset, which focuses on the production of riboflavin (vitamin B2) from various Bacillus subtilis. This dataset, provided by the Dutch State Mines Nutritional Products, is accessible via hdi R package (Dezeure et al., 2015). It comprises a single continuous response variable that represents the logarithm of the riboflavin production rate, alongside $p=4,088$ covariates, which correspond to the logarithm of expression levels for 4,088 bacterial genes. The primary objective of analyzing this dataset is to identify genes that influence riboflavin production, with the ultimate goal of genetically engineering bacteria to enhance riboflavin yield. Data were collected from $n=71$ relatively homogeneous samples, which were repeatedly hybridized during a feed-batch fermentation process involving different engineered strains and varying fermentation conditions. Bühlmann et al. (2014) employed stability selection with Lasso and identified three genes—LYSC_at, YOAB_at, and YXLD_at—as relevant genes in this problem.
@@ 4 Results
In this section, we illustrate the efficacy of the proposed methodology through its application to the synthetic and real datasets introduced in Section 3. We used the GitHub repository associated with Nogueira et al. (2018) to implement their stability estimator and to obtain confidence intervals [^1] https://github.com/nogueirs/JMLR2018.
@@
Figure 1 illustrates the selection stability of the stability selection results on a grid of regularization values, applied to the synthetic datasets introduced in Section 3, using three different values of $\rho$ and choosing the number of sub-samples $B=500$. The horizontal dashed red lines indicate the stability thresholds of $0.4$ and $0.75$ as discussed in Section 2. As illustrated in Figure 1, the regularization value that minimizes the cross-validation error, $\lambda_{\text{min}}$, falls within the poorly stable region. In addition, the regularization value that maintains the cross-validation error within one standard error of $\lambda_{\text{min}}$, referred to as $\lambda_{\text{1se}}$ (Hastie et al., 2009), lies within the poor to intermediate stability regions. In contrast, $\lambda_{\text{stable}}$, introduced in Equation (3), is specifically designed to fall within the region of excellent stability, where applicable.
@@
To demonstrate the convergence of the stability estimator $\hat{\Phi}(\cdot)$, for each correlation scenario, we used $\lambda_{\text{stable}}$ and calculated the stability of the selection matrix $M(\lambda_{\text{stable}})$ through sequential sub-sampling, using Equation (2). The stability values across iterations are shown in Figure 3. The blue shading around the line represents the $95\%$ confidence interval for $\hat{\Phi}(\cdot)$. As can be observed, the interval narrows with increasing iterations, indicating a reduction in the estimator’s uncertainty. Figure 3 reveals that, across all three scenarios, after approximately $B^{*}\approx 200$ iterations, the stability estimator converges, with no significant change thereafter in its value. By monitoring the stability trajectory throughout the process, valuable insight can be gained in determining the optimal cut-off point, that is, the optimal number of sub-samples in terms of stability of the results.
@@
Next, we apply stable stability selection to the riboflavin dataset. As described in Section 3, the dataset consists of $p=4,088$ genes, the main objective being to identify the key genes that influence riboflavin production. As before, we choose the number of sub-samples $B=500$. The predictor variables are standardized using the scale function prior to being input into Lasso.
In this problem, $\lambda_{\text{stable}}$ does not exist, so we use $\lambda_{\text{stable-1sd}}$ instead. After $B$ iterations, four genes YXLD_at ($\hat{\Pi}_{\texttt{YXLD\_at}}^{\lambda_{\text{stable-1sd}}}=0.606$), YOAB_at ($\hat{\Pi}_{\texttt{YOAB\_at}}^{\lambda_{\text{stable-1sd}}}=0.558$), LYSC_at ($\hat{\Pi}_{\texttt{LYSC\_at}}^{\lambda_{\text{stable-1sd}}}=0.540$), and YCKE_at ($\hat{\Pi}_{\texttt{YCKE\_at}}^{\lambda_{\text{stable-1sd}}}=0.532$) have selection frequencies greater than $0.5$.
Figure 4 illustrates the stability of stability selection when applied to the riboflavin dataset. As shown in Figure 4(a), $\lambda_{\text{stable}}$ does not exist in this example, as the stability values do not surpass the $0.75$ threshold. Figure 4(b), generated with $\lambda_{\text{stable-1sd}}$, indicates the convergence slightly above $0.2$ after about $200$ iterations. These results imply that stability selection with Lasso exhibits poor selection stability on this dataset.
@@
Finally, we apply the proposed methodology to the rat microarray data. As mentioned in Section 3, the data consists of $p=3,083$ gene probes. The main aim for this data is to identify probes that influence the TRIM32 probe. Again, we choose the number of sub-samples $B=500$. The predictor variables are standardized using the scale function prior to being input into Lasso.
Due to the instability of the results, $\lambda_{\text{stable}}$ does not exist in this problem. After $B$ iterations, three probes have selection frequencies greater than $0.5$: probe 1390539_at ($\hat{\Pi}_{\texttt{1390539\_at}}^{\lambda_{\text{stable-1sd}}}=0.640$), probe 1389457_at ($\hat{\Pi}_{\texttt{1389457\_at}}^{\lambda_{\text{stable-1sd}}}=0.570$), and probe 1376747_at ($\hat{\Pi}_{\texttt{1376747\_at}}^{\lambda_{\text{stable-1sd}}}=0.564$).
As shown in Figure 5, similar to the results of the riboflavin dataset, the procedure cannot be considered stable. Specifically, Figure 5(a) reveals that $\lambda_{\text{stable}}$ does not exist, as the stability values do not exceed the threshold of $0.75$. For Figure 5(b), we used $\lambda_{\text{stable-1sd}}$. Figure 5(b) shows that the stability values converged to approximately $0.15$.
@@
The convergence plots for both synthetic and real examples indicate that, regardless of whether the convergence value is high or low, stability values tend to stabilize after $B^{*}\approx 200$ iterations. This observation provides a useful rule of thumb for selecting the required number of sub-samples in the stability selection framework when using the Lasso.
@@
The two assumptions in Corollary 1 are both justifiable. Increasing the regularization value is expected to allow the model to achieve an optimal balance of sparsity by shrinking irrelevant variables, thereby enhancing stability. The regularization value $\lambda_{\text{stable}}$ is intended to signify the point where the model attains high stability with minimal loss of accuracy. Consequently, we anticipate that the stability curve exhibits a non-decreasing trend prior to reaching $\lambda_{\text{stable}}$.
Regarding the second assumption, it is expected that increasing regularization helps the model eliminate irrelevant variables, improving its ability to predict the response variable. The regularization values $\lambda_{\text{min}}$ and $\lambda_{\text{1se}}$ are designed to capture this behavior, marking the points where regularization is most effective in terms of predictive ability. Beyond these points, we anticipate that the loss curve will be non-decreasing. Since $\lambda_{\text{stable}}$ is intended to trade a small amount of accuracy for greater stability, we expect $\lambda_{\text{stable}}$ to occur after $\lambda_{\text{1se}}$, that is, $\lambda_{\text{stable}}>\lambda_{\text{1se}}>\lambda_{\text{min}}$, which implies that the loss curve is expected to be non-decreasing after $\lambda_{\text{stable}}$.

### v3

@@ Abstract
Stability selection is a widely adopted resampling-based framework for high-dimensional variable selection. This paper seeks to broaden the use of an established stability estimator to evaluate the overall stability of the stability selection results, moving beyond single-variable analysis. We suggest that the stability estimator offers two advantages: it can serve as a reference to reflect the robustness of the results obtained, and it can help identify a Pareto optimal regularization value to improve stability. By determining the regularization value, we calibrate key stability selection parameters, namely, the decision-making threshold and the expected number of falsely selected variables, within established theoretical bounds. In addition, the convergence of stability values over successive sub-samples sheds light on the required number of sub-samples addressing a notable gap in prior studies. The stabplot R package is developed to facilitate the use of the methodology featured in this paper.
Keywords: Bioinformatics, Feature Selection, Hyperparameter Tuning, Lasso, Regularization Tuning, Variable Selection
@@ 1 Introduction
Stability selection, a widely recognized resampling-based variable selection framework, performs selection by examining how frequently variables are chosen across multiple randomly selected resamples (Meinshausen and Bühlmann, 2010). Meinshausen and Bühlmann (2010) analyzed the selection frequencies for individual variables over a range of regularization values, presenting the results in a plot known as the ‘stability path’. This plot facilitates the identification of variables that are consistently selected over the regularization grid. In this paper, we shift the focus from examining the stability of individual variables to evaluating the overall stability of the stability selection results. Employing an established stability estimator, we assess the stability of the results in its entirety and introduce what we call ‘stable stability selection’, which identifies the variables that are associated with highly stable results.
@@
The stability of a variable selection method refers to the consistency with which it selects variables across different training sets drawn from the same underlying distribution (Kalousis et al., 2007). The stability selection approach described by Meinshausen and Bühlmann (2010) uses sub-samples that are half the size of the sample size of the original dataset to perform variable selection. Stability selection enables the identification of variables that are consistently selected as relevant across the majority of sub-samples given a regularization value, considering them as stable variables. Furthermore, Meinshausen and Bühlmann (2010) established an asymptotic upper-bound for the Per-Family Error Rate, which represents the expected number of falsely selected variables. Later, Shah and Samworth (2013) introduced the complementary pairs stability selection method, which computes selection frequencies by counting how often a variable is included in models fitted to complementary $50\%$ sub-samples. Shah and Samworth (2013) introduced stricter upper-bounds for the expected number of variables selected with low selection frequencies.
@@
There is a substantial body of research on the stability of variable selection methods, where numerous stability measures have been proposed. For more comprehensive information, we refer to Kuncheva (2007), Nogueira et al. (2018), and Sen et al. (2021). Nogueira et al. (2018) conducted an extensive literature review on the topic, consolidating the desirable properties of stability measures into five mathematical conditions. Nogueira et al. (2018) showed that none of the stability measures previously introduced satisfied all five conditions. In response, Nogueira et al. (2018) proposed a novel stability measure, which was the first in the literature to meet all these criteria, generalizing several previous works such as in Kuncheva (2007).
@@
In this paper, we apply the stability estimator proposed by Nogueira et al. (2018) to the stability selection results, and proxify the stability values to find the optimal regularization value that achieves high stability with the least possible loss in prediction accuracy. The optimal regularization value then is employed to calibrate two critical parameters of stability selection: the decision-making threshold and the expected number of falsely selected variables, balancing them with respect to one another. We also use the optimal regularization value to define a new variable selection criterion, called stable stability selection, prioritizing the selection of variables associated with highly stable results, rather than those mostly selected in their best-case scenarios over the regularization grid, which is the primary criterion in the traditional stability selection framework. In addition, the convergence value of the stability estimator over the successive sub-samples serves as a reference to assess the reliability of the results obtained. The point at which convergence occurs provides valuable insight into the required number of sub-samples, an aspect for which we found no dedicated research addressing its determination.
@@
Nogueira et al. (2018) suggested that stability and prediction accuracy can be interpreted through the lens of a Pareto front Pareto (1896). We refer to this methodology as stability-accuracy selection. Building on this perspective, we demonstrate that, given two justifiable assumptions, our regularization value constitutes a Pareto-optimal solution within this context.
@@
Hyper-parameter tuning can be approached from various perspectives (Aftab et al., 2025). The use of stability to enhance variable selection has been a subject of interest since at least the work of Liu et al. (2010). Among the existing literature, the study by Sun et al. (2013) is perhaps the most closely related to the present work and has continued to attract interest, as demonstrated by contributions such as Yang et al. (2020), Reavie (2021), Wen et al. (2023), and Lee and Chen (2024). In Sun et al. (2013), the regularization parameter of penalized regression models is selected based on a stability-driven criterion. Nonetheless, our methodology departs from that of Sun et al. (2013) by being fully embedded within the stability selection framework, using its internal selection outcomes to guide regularization tuning. In contrast, their method is designed to function externally, resulting in greater computational overhead due to its decoupled implementation.
@@
The rest of this paper is organized as follows. Section 2 outlines the proposed methodology. Section 3 describes the real and synthetic datasets used in the paper. The numerical results of applying the proposed method, along with the demonstration of Pareto optimality of the proposed regularization value, are presented in Section 4.
@@ 2 Methodology
The following procedure is applied to each regularization value $\lambda\in\Lambda$. In each iteration of the stability selection process, a sub-sample, comprising half the size of the original dataset $\mathcal{D}$, is randomly selected. The Lasso model is then fitted to this sub-sample using the current value of $\lambda$, and the binary selection outcomes are recorded as a row in the binary selection matrix $M(\lambda)\in\operatorname{mat}(B\times p)$, where $B$ denotes the number of sub-samples. Therefore, in the end, we obtain $|\Lambda|$ distinct selection matrices $M(\lambda)$. In this context, $M(\lambda)_{bj}=1$ indicates that the $j$th variable is identified as part of the signal set $\hat{S}(\lambda)$ when Lasso is applied to the $b$th sub-sample given $\lambda$. In contrast, $M(\lambda)_{bj}=0$ signifies that the $j$th variable is classified as belonging to the noise set $\hat{N}(\lambda)$ after applying Lasso to the $b$th sub-sample given $\lambda$.
@@
where $s_{j}^{2}$ denotes the unbiased sample variance of the binary selection statuses of the $j$th variable, while $q(\lambda)$ denotes the average number of variables selected under the regularization parameter $\lambda$. The stability estimator $\hat{\Phi}(\cdot)$ is bounded by $\left[-\frac{1}{B-1},1\right]$ (Nogueira et al., 2018); the larger the $\hat{\Phi}(M(\lambda))$, the more stable is $M(\lambda)$. To interpret the stability values of Equation (2), Nogueira et al. (2018) adopted the guidelines proposed by Fleiss et al. (2004). According to Nogueira et al. (2018), stability values that exceed $0.75$ indicate excellent agreement between selections beyond what would be expected by random chance, while values below $0.4$ signify poor agreement between them. Based on Nogueira et al. (2018), stability values that reside between these thresholds are categorized as indicative of intermediate to good stability.
@@
The threshold value of $0.75$ is somewhat arbitrary and may not be attainable in certain situations. As we will see in Section 4, $\lambda_{\text{stable}}$ may not exist in certain practical applications due to instability of the results, which prevents the stability values from exceeding $0.75$. In such cases, aligning with the $\lambda_{\text{1se}}$ strategy introduced by Hastie et al. (2009), we propose
@@
Both upper-bounds of the Per-Family Error Rate, mentioned in Section 1, provided by Meinshausen and Bühlmann (2010) and Shah and Samworth (2013), depend on the values of $\lambda$ and $\pi_{\text{thr}}$. The calibration of these two values requires the arbitrary selection of one of these two parameters, which can be challenging to justify (Bodinier et al., 2023). By adopting $\lambda=\lambda_{\text{stable}}$ or $\lambda=\lambda_{\text{stable-1sd}}$, the upper-bound can be explicitly tailored to the data owner’s preferences. If a pre-determined $\pi_{\text{thr}}$ is chosen, the Per-Family Error Rate is deterministically obtained; alternatively, if a fixed pre-determined Per-Family Error Rate value is desired, the corresponding $\pi_{\text{thr}}$ is enforced.
@@
As outlined in Section 1, as the number of sub-samples increases, the convergence of $\hat{\Phi}(\cdot)$ to the population stability $\Phi$ is ensured. To illustrate this, we define $M^{(t)}(\lambda)$ as the matrix containing the selection results from the first $t$ sub-samples for a given $\lambda\in\Lambda$, that is, the first $t$ rows of $M(\lambda)$. The objective is to evaluate the stability of $M^{(t)}(\lambda)$ across successive sub-samples in order to determine the appropriate cut-off point for the process, that is, the number of sub-samples required to achieve convergence in stability values. Beyond this threshold, additional sub-samples do not significantly change the stability of $M(\lambda)$. We propose plotting the stability values of the selection matrix over the sequential sub-sampling, that is $\hat{\Phi}(M^{(t)}(\lambda));\;t=2,3,\ldots B$, against $t$ to monitor the convergence status of the stability values. Given that the asymptotic distribution of $\hat{\Phi}(\cdot)$ is established, a confidence bound can be drawn along the curve to reflect the inherent uncertainty of the stability estimator.
In Section 4, we demonstrate the Pareto optimality of $\lambda_{\text{stable}}$, and present the numerical results obtained from the application of the proposed approach to synthetic and real datasets, which are introduced in Section 3.
@@ 3 Datasets
We evaluate our methodology by performing evaluations on synthetic and two real bioinformatics datasets, as detailed below.
@@
As synthetic data, we consider a dataset with a sample size of $n=50$ that includes $p=500$ predictor variables. The predictor variables $\boldsymbol{x}^{\intercal}_{i}$ are independently drawn from the Normal distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma\in\operatorname{mat}(p,p)$ has a diagonal of ones and $\sigma_{jk}=\rho^{|j-k|}$ for $j\neq k$ where $\rho\in\{0.2,0.5,0.8\}$. The response variable is linearly dependent solely on the first two predictor variables, characterized by the coefficient vector $\boldsymbol{\beta}=(1.5,1.1,0,\ldots 0)^{\top}$, and the error term $\boldsymbol{\varepsilon}$ is an i.i.d. sample from the standard Normal distribution $\mathcal{N}(0,1)$.
@@
For the first real example, we use the well-established ‘Riboflavin’ dataset, which focuses on the production of riboflavin (vitamin B2) from various Bacillus subtilis. This dataset, provided by the Dutch State Mines Nutritional Products, is accessible via hdi R package (Dezeure et al., 2015). It comprises a single continuous response variable that represents the logarithm of the riboflavin production rate, alongside $p=4,088$ covariates, which correspond to the logarithm of expression levels for 4,088 bacterial genes. The primary objective of analyzing this dataset is to identify genes that are associated with riboflavin production, with the ultimate goal of genetically engineering bacteria to enhance riboflavin yield. Data were collected from $n=71$ relatively homogeneous samples, which were repeatedly hybridized during a feed-batch fermentation process involving different engineered strains and varying fermentation conditions. Bühlmann et al. (2014) employed stability selection with Lasso and identified three genes—LYSC_at, YOAB_at, and YXLD_at—as relevant genes in this problem.
@@ 4 Results
In this section, we illustrate the efficacy of our methodology through its application to the synthetic and real datasets introduced in Section 3. We also demonstrate in Corollary 1 that $\lambda_{\text{stable}}$ constitutes a Pareto optimal solution in the trade-off between stability and accuracy in variable selection. We used the GitHub repository associated with Nogueira et al. (2018) to implement their stability estimator and to obtain confidence intervals [^1] https://github.com/nogueirs/JMLR2018.
@@
Figure 1 illustrates the selection stability of the stability selection results on a grid of regularization values, applied to the synthetic datasets introduced in Section 3, using three different values of $\rho$ and choosing the number of sub-samples $B=500$. The horizontal dashed red lines indicate the stability thresholds of $0.4$ and $0.75$ as discussed in Section 2. As illustrated in Figure 1, the regularization value that minimizes the cross-validation error, $\lambda_{\text{min}}$, falls within the poorly stable region. In addition, $\lambda_{\text{1se}}$ lies within the poor to intermediate stability regions. In contrast, $\lambda_{\text{stable}}$, introduced in Equation (3), is specifically designed to fall within the region of excellent stability, where applicable.
@@
To demonstrate the convergence of the stability estimator $\hat{\Phi}(\cdot)$, for each correlation scenario, we used $\lambda_{\text{stable}}$ and calculated the stability of the selection matrix $M(\lambda_{\text{stable}})$ through sequential sub-sampling, using Equation (2). The stability values across iterations are shown in Figure 3. The blue shading around the line represents the $95\%$ confidence interval for $\hat{\Phi}(M(\lambda_{\text{stable}}))$. As can be observed, the interval narrows with increasing iterations, indicating a reduction in the estimator’s uncertainty. Figure 3 reveals that, across all three scenarios, after approximately $B^{*}\approx 200$ iterations, the stability estimator converges, with no significant change thereafter in its value. By monitoring the stability trajectory throughout the process, valuable insight can be gained in determining the optimal cut-off point, that is, the optimal number of sub-samples in terms of stability of the results.
@@
Next, we apply stable stability selection to the riboflavin dataset. As described in Section 3, the dataset consists of $p=4,088$ genes, the main objective being to identify the key genes that are associated with riboflavin production. As before, we choose the number of sub-samples $B=500$. The predictor variables are standardized using the scale function prior to being input into Lasso. In this problem, $\lambda_{\text{stable}}$ does not exist, so we use $\lambda_{\text{stable-1sd}}$ instead. After $B$ iterations, four genes YXLD_at ($\hat{\Pi}_{\texttt{YXLD\_at}}^{\lambda_{\text{stable-1sd}}}=0.606$), YOAB_at ($\hat{\Pi}_{\texttt{YOAB\_at}}^{\lambda_{\text{stable-1sd}}}=0.558$), LYSC_at ($\hat{\Pi}_{\texttt{LYSC\_at}}^{\lambda_{\text{stable-1sd}}}=0.540$), and YCKE_at ($\hat{\Pi}_{\texttt{YCKE\_at}}^{\lambda_{\text{stable-1sd}}}=0.532$) have selection frequencies greater than $0.5$.
Figure 4 illustrates the stability of stability selection results when applied to the riboflavin dataset. As shown in Figure 4(a), $\lambda_{\text{stable}}$ does not exist in this example, as the stability values do not surpass the $0.75$ threshold. Figure 4(b), generated with $\lambda_{\text{stable-1sd}}$, indicates the convergence slightly above $0.2$ after about $200$ iterations. These results imply that stability selection with Lasso exhibits poor selection stability on this dataset.
@@
Finally, we apply our methodology to the rat microarray data. As mentioned in Section 3, the data consists of $p=3,083$ gene probes. The main aim for this data is to identify probes that are associated with the TRIM32 probe. Again, we choose the number of sub-samples $B=500$. The predictor variables are standardized using the scale function prior to being input into Lasso. Due to the instability of the results, $\lambda_{\text{stable}}$ does not exist in this problem. After $B$ iterations, three probes have selection frequencies greater than $0.5$: probe 1390539_at ($\hat{\Pi}_{\texttt{1390539\_at}}^{\lambda_{\text{stable-1sd}}}=0.640$), probe 1389457_at ($\hat{\Pi}_{\texttt{1389457\_at}}^{\lambda_{\text{stable-1sd}}}=0.570$), and probe 1376747_at ($\hat{\Pi}_{\texttt{1376747\_at}}^{\lambda_{\text{stable-1sd}}}=0.564$).
As shown in Figure 5, similar to the results of the riboflavin dataset, the results cannot be considered stable. Specifically, Figure 5(a) reveals that $\lambda_{\text{stable}}$ does not exist, as the stability values do not exceed the threshold of $0.75$. For Figure 5(b), we used $\lambda_{\text{stable-1sd}}$. Figure 5(b) shows that the stability values converged to approximately $0.15$.
@@
The two assumptions in Corollary 1 are both justifiable. Increasing the regularization value from zero is expected to allow the model to achieve an optimal balance of sparsity by shrinking irrelevant variables, thereby enhancing stability. The regularization value $\lambda_{\text{stable}}$ is intended to signify the point where the model attains high stability with minimal loss of accuracy. Consequently, we anticipate that the stability curve exhibits a non-decreasing trend prior to reaching $\lambda_{\text{stable}}$.
Regarding the second assumption, it is expected that increasing regularization helps the model eliminate irrelevant variables, improving its ability to predict the response variable. The regularization values $\lambda_{\text{min}}$ and $\lambda_{\text{1se}}$ are designed to capture this behavior, marking the points where regularization is most effective in terms of predictive ability. Beyond these points, we anticipate that the loss curve will be non-decreasing as a consequence of model over-shrinkage. Since $\lambda_{\text{stable}}$ is intended to trade a small amount of accuracy for greater stability, we expect $\lambda_{\text{stable}}$ to occur after $\lambda_{\text{1se}}$, that is, $\lambda_{\text{stable}}>\lambda_{\text{1se}}>\lambda_{\text{min}}$, which implies that the loss curve is expected to be non-decreasing after $\lambda_{\text{stable}}$.
@@
Although the methodology is presented within the context of the Lasso, its extension to other regularization-based variable selection techniques remains a natural and conceptually supported direction, as suggested by Sun et al. (2013).